OpenStudy (adajiamcneal):

Will Medal :) y=-3(1/3)^2+2(1/3)+5 parenthesis are fractions.

OpenStudy (anonymous):

y=163

OpenStudy (adajiamcneal):

how did you get it

OpenStudy (anonymous):

(16/3)

OpenStudy (anonymous):

http://www.cymath.com/answer.php

OpenStudy (adajiamcneal):

gracias :)

OpenStudy (anonymous):

denada mija ;)

jimthompson5910 (jim_thompson5910):

\[\Large y = -3\left(\frac{1}{3}\right)^2 + 2\left(\frac{1}{3}\right)+5\] \[\Large y = -3\left(\frac{1}{9}\right) + 2\left(\frac{1}{3}\right)+5\] \[\Large y = \frac{-3}{9} + \frac{2}{3}+5\] \[\Large y = \frac{-1}{3} + \frac{2}{3}+5\] \[\Large y = \frac{-1}{3} + \frac{2}{3} + \frac{15}{3}\] \[\Large y = \frac{-1+2+15}{3}\] \[\Large y = \frac{16}{3}\]

OpenStudy (adajiamcneal):

jim he helped me already pero thanks for trying :)

jimthompson5910 (jim_thompson5910):

well it helps to know why/how the answer is 16/3

OpenStudy (adajiamcneal):

yeah I know :) he gave me a website that showed me pero ill give you a medal on some one else's question that you answered haha